OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence?5, 25, 125, 250, …

OpenStudy (anonymous):

please help me someone

OpenStudy (unklerhaukus):

what factor does the next term in the sequence have compared the the term before it

OpenStudy (anonymous):

its times 5

OpenStudy (unklerhaukus):

good ,

OpenStudy (anonymous):

so what would be the formula

OpenStudy (unklerhaukus):

so the first term = 5 the second is 5x5 = 5^2 the third is 5x5x5 = 5^3

OpenStudy (unklerhaukus):

whats the nth term?

OpenStudy (anonymous):

i not sure

OpenStudy (unklerhaukus):

well when n=1 the term is 5 when n=2 the term is 5^2 when n=3 the term is 5^3 when n=n the term is ......?

OpenStudy (anonymous):

5^9

OpenStudy (anonymous):

so is the formula an=5(5)^n − 1

OpenStudy (unklerhaukus):

why do you have two fives in you equation?

OpenStudy (unklerhaukus):

oh wait i see now

OpenStudy (unklerhaukus):

yeah you could have that

OpenStudy (anonymous):

ya

OpenStudy (unklerhaukus):

\[a_n=5(5)^{n − 1}\] \[a_1=5(5)^{1 − 1}=5(5)^0=5\times1=5\]\[a_2=5(5)^{2 − 1}=5(5)^1=5\times5=25\] it works

OpenStudy (unklerhaukus):

but you could simplify it \[a_n=5(5)^{n − 1}=5^1\times5^{n-1}=5^{n-1+1}=5^n\]